OpenStudy (anonymous):

Using Pythagorean Theorem- A plane, intending to fly 252 miles due west from town A to town B, drifts off course and ends up 39 miles due south of town B. How far did the plane fly in getting to town B?

OpenStudy (anonymous):



OpenStudy (anonymous):

Solve for c, does that help?

OpenStudy (anonymous):

Yep! Thank you!

OpenStudy (crystal_bliss):

\[a^2+b^2=c^2\]

OpenStudy (crystal_bliss):

so what did you get for your answer?

OpenStudy (anonymous):

Giving that A is 39 and B is 252 39^2 + 252^2 = √65025 which can simplify to 255 miles. The picture really helped.

OpenStudy (anonymous):

As far as I'm concerned that should be right.

OpenStudy (anonymous):

Yup thats right